OpenStudy (calculusxy):

Richard and Kayla have 48 apples. Kayla has 35% more apples than Richard. a) Write an equation that represents this situation.

OpenStudy (calculusxy):

@confluxepic

OpenStudy (amorfide):

okay if I had 10 and I increased it by 35% I would have to find 35% of 10 then add it on so 10 is 100% to get 35% of 10 I do \[\frac{ 35 \times 10 }{ 100 }\] \[\frac{ 35 }{ 10 }=3.5\] a 35% increase of 10 would be 10+ 3.5 this is 100% + 35% this in total is 135% 135% as a decimal is 1.35 if you let apples be the variable a we know that kayla has 35% more apples than richard this means 135% of richards apples, is the amount that kayla has so let r= amount of apples richard has and let k=kayla we get an equation that kayla is equal to 135% of richard so we get k=1.35r

OpenStudy (calculusxy):

@amorfide Thank you.

OpenStudy (confluxepic):

Well done. @amorfide

OpenStudy (amorfide):

thank you :)